OpenStudy (anonymous):

limit of f(x) given by f(x)=(1-cosx)/x^2*cos4x, x->0

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

great, now i dont want to give you my answer cuase its prolly wrong

OpenStudy (nikvist):

\[\lim_{x\rightarrow 0}\frac{1-\cos{x}}{x^2\cdot\cos{4x}}= \underbrace{\lim_{x\rightarrow 0}\frac{1}{\cos{4x}}}_{=1}\cdot\lim_{x\rightarrow 0}\frac{1-\cos{x}}{x^2}=\lim_{x\rightarrow 0}\frac{1-\cos{x}}{x^2}=\] \[=\lim_{x\rightarrow 0}\frac{\sin{x}}{2x}=\lim_{x\rightarrow 0}\frac{\cos{x}}{2}=\frac{1}{2}\]